How many grams of solute are present in 845ml of 0.670 M KBr?
1
Expert's answer
2019-12-03T08:26:48-0500

"c=\frac{n}{V}"

"n=c\times V"


"n(KBr) = 0.670\times0.845 = 0.5662 mol"


"m(KBr)= M\times n = 119.00 \times 0.5662 = 67.37 g"